TMS Meeting Minutes, January 21, 2021

Attendees: Andrew, Breann, Heejoo, John, Trenton, Olga, Yang (Zoom Meeting)
  • Rain in Tucson, snow at the LBT
  • Heejoo visited the mountain
    • Mostly maintenance on DX side, cleaning of collimators
    • Fiber routing on SX side, 12 fiber channels available (9 channels for LBC, 2 channels for diameter, 1 extra)
    • Unfortunately, one fiber is broken, so only 11 ch working
    • Had a quick meeting with Mike Gardener after handing over the new retro
    • He is planning to visit the LBT next week or later
    • Planning to continue work on fiber routing to mirror cell
    • Heejoo will return later to place the fibers in the proper location once collimators are installed
    • Current eTalon MPO casette does not give a reasonable signal
    • Unfortunately, the replacement Heejoo bought is the wrong model
    • Heejoo reached out to several manufacturers to find the right model for our system
    • The problem seems to lie with the connection from multi-core fibers into the MPO casette
  • Andrew suggests sending the current MPO casette back to eTalon for replacement
  • The fiber switch (stored in storage room, unused) for toggling between SX and DX appears to only have two inputs, a LAN input and power (seems we need a specific controller for this hardware)
  • Mattheiu has completed the TMS Laser Alarm email alert system, and it has worked perfectly over the past two weeks
  • Two days ago, software crash error
    • Laser was on, software crashed, shutter opened, alarm closed the shutter again
    • Corresponding with eTalon engineer Mark
    • At the time of failure, the TCP/IP server was NOT connected
    • Interesting contradiction, that the shutter opened when the software crashed (opposite is true for open; if shutter open, stays open upon crash)
    • eTalon working on stopping the software from crashing
    • To be safe, we should probably keep the laser turned off until we can solve this problem (expend the 20,000 cycle lifetime)
    • Heejoo proposes implementing a system of checking for TMS on/off when conducting LBC observations
    • Andrew notes that the best approach wouldn't involve any human interventention
    • Once we are convinced that the laser alarm works, we should rely on this to control the undesired states
    • Heejoo notes that the laser alarm references DMS data to readout laser power from ch 1 of TMS; alarm sent out based on voltage level; alarms only on transitions between states
  • Collimator Search
    • Want a collimator optimized over 1.5 um to 0.6 um
    • Heejoo suggests we can find a similar fiber collimator from Edmund Optics or ThorLabs
    • Has been conducting some reverse engineering on the large fiber collimators currently in use on the mountain
    • Suggests that we may be able to make a simple adjustment, such as the fiber tip position relative to the collimator, to have reasonable performance over both wavelengths
    • Trenton working on finding a solution, as the SX install is approaching quickly
    • Heejoo plans to test the proposed collimator on the mountain once we have some samples in hand and pick the best to outfit the SX mirror with
    • Heejoo ordered some of the other hardware necessary to set up the fiber collimators on the mirror cell
    • Andrew suggests using an NZ company, Kiwi Optics, to manufacture a couple hundred collimators
  • Andrew notes that alignment generally looks great
  • Heejoo noticed that most of the fibers had deviated 3-4 mm (when observing the returned spot) for every channel, so the realignment was the main driving factor in the channel strength improvement
  • Cleaning was also very important for certain collimators, where dust was most present
  • Lots of closed dome time in the coming days
  • Would like to run the system through its paces
  • Of note, coordination usually managed over the polycom channel
  • Andrew would like to see signal strength over the elevation range to test improvement with cleaned hardware
  • Olga asked about hub rotation with elevation changes on LBCr
    • We are currently using a model (collimation lookup table) to provide hub rotation for a coma-neutral rotation over elevation ranges
    • Lookup table referenced to correct the hub for elevation change
    • Twist of the hub with elevation has to be controlled with the collimation table
    • Around Zenith, a sort of rocking observed, not very consistent nor repeatable
    • Consider commanding 0.5 deg steps near Zenith to probe this (consider range around 3 deg from Zenith)
    • Strange, as you would expect the jump to happen around 30 deg when the gravity loading starts to let up, but it happens when everything is under load
  • Wide Field Retro Testing (HeNe 633 nm)
    • Heejoo tested Andrew's retro
    • At near normal, 1.9 microW return
    • At ~50 deg incidence angle, getting ### microW return
      • The optimizer was designed for 1550 nm, not HeNe, so you get substantial Fresnel reflections off the front
    • Andrew asks that Heejoo work with Mike to set up a test to use three ball bearings for a back surface mount to set up the solid retro as a sort of SMR
      • Interested in OPD stability as the retro is rotated on the ball bearings and the signal gain
    • Andrew would also like Heejoo to look at the solid retro with a Leica IR laser tracker (Mike has it at the moment)
  • Yang working on implementing the StdDev based channel filter
  • Andrew planning to send out his Mathematica file with the implementation this weekend

-- TrentonBrendel - 21 Jan 2021
Topic revision: r2 - 22 Jan 2021, TrentonBrendel
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback